OpenStudy (anonymous):

A parallelogram with vertices (–4, 1), (1, 3), (2, 0), and (–3, –2) is translated left 3 units, up 5 units. What are the coordinates of the vertices of the translated parallelogram? A. (–1, 6), (4, 8), (5, 5), and (0, 3) B. (1, –2), (6, 0), (7, –3), and (2, –5) C. (–7, 6), (–2, 5), (–1, 5), and (–6, 3) D. (–7, 6), (–2, 8), (–1, 5), and (–6, 3)

OpenStudy (gabbyalicorn):

do you have like a picture?

OpenStudy (anonymous):

No, they didn't give a graph, which stinks

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Im here to help. one sec

OpenStudy (welshfella):

work out the new coordinates by subtracting / adding as follows take the point (-4,1) left 3 units make -4 go to -4-3 = -7 up 5 units 1 goes to 6 so new coordinates are (-7,6) which means the answer is either c or d

OpenStudy (welshfella):

do same process with the next point (1 , 3) etc

OpenStudy (welshfella):

@Music_me - do you follow what I'm saying?

OpenStudy (anonymous):

Yes, I'm just trying them out

OpenStudy (welshfella):

good

OpenStudy (anonymous):

i would help.. but @welshfella cant handle anything true

OpenStudy (welshfella):

Point (1,3);- 1-3 = ? 3 + 5 = ?

OpenStudy (anonymous):

(-2,8) rite?

OpenStudy (welshfella):

right on

OpenStudy (anonymous):

Okie

OpenStudy (welshfella):

so now you should know the correct option...

OpenStudy (anonymous):

Would it be D?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

Okie, thnx! :)

OpenStudy (welshfella):

yw